# LiT: Delving into a Simplified Linear Diffusion Transformer for Image Generation

### v1

@@ Abstract
In commonly used sub-quadratic complexity modules, linear attention benefits from simplicity and high parallelism, making it promising for image synthesis tasks. However, the architectural design and learning strategy for linear attention remain underexplored in this field. In this paper, we offer a suite of ready-to-use solutions for efficient linear diffusion Transformers. Our core contributions include: (1) Simplified Linear Attention using few heads, observing the free-lunch effect of performance without latency increase. (2) Weight inheritance from a fully pre-trained diffusion Transformer: initializing linear Transformer using pre-trained diffusion Transformer and loading all parameters except for those related to linear attention. (3) Hybrid knowledge distillation objective: using a pre-trained diffusion Transformer to help the training of the student linear Transformer, supervising not only the predicted noise but also the variance of the reverse diffusion process. These guidelines lead to our proposed Linear Diffusion Transformer (LiT), an efficient text-to-image Transformer that can be deployed offline on a laptop. Experiments show that in class-conditional 256256 and 512512 ImageNet benchmark LiT achieves highly competitive FID while reducing training steps by and compared to DiT. LiT also rivals methods based on Mamba or Gated Linear Attention. Besides, for text-to-image generation, LiT allows for the rapid synthesis of up to 1K resolution photorealistic images. Project page: https://techmonsterwang.github.io/LiT/.
@@ 1 Introduction
Diffusion models [29, 67] based on Transformer [71], such as PixArt- [7] and Stable Diffusion 3 [16], recently demonstrate potential commercial value in the field of high-resolution text-to-image generation. However, one of the core computational modules i.e., self-attention, exhibits a quadratic complexity with respect to sequence length, leading to increased latency and GPU memory usage in high-resolution scenarios, which makes it less feasible for resource-limited edge devices, as illustrated in Fig. 4 and Fig. 8 (Sec. B of the appendix).
Accordingly, existing works [69, 87] adopt sub-quadratic modules, such as State Space Models (SSM) [20, 21] and Gated Linear Attention (GLA) [78], to construct diffusion models for image generation. Their designs are illustrated in Fig. 2-(a). While they all achieve sub-quadratic complexity, the overall complexity differs substantially. First, both SSM blocks and GLA rely on recurrent states, leading to iterative computation that limits parallelization opportunities. Second, these modules have complex computational graphs and involve low-arithmetic-intensity operations (e.g., element-wise multiplication).
In contrast, linear attention [38, 3, 23, 1, 6] linear attention provides an efficient alternative to self-attention, with a simple design and easy parallelization. It can be observed that linear attention speeds up generation compared to self-attention at the same resolution. For example, at a 2048 resolution, linear attention is nearly 9 faster (shown in Fig. 4-(a) and (b)). Besides, linear attention allows for fewer GPU memory for higher-resolution image generation. For instance, the GPU memory for generating 2048px image with DiT-S/2 can be dropped from 14GB to 4GB if replacing with the linear attention counterpart. Therefore, training a high-performance diffusion Transformer with linear attention is valuable for edge-side devices with strict latency and computational resource requirements.
Nevertheless, how to effectively and rapidly train linear attention-based diffusion Transformer for challenging image generation tasks? On one hand, although linear attention successfully replicates the effectiveness of self-attention in visual recognition, architectural designs for linear attention in image generation remain an under-explored issue. On the other hand, training diffusion models from scratch is costly. For example, as statistically reported in [8], training RAPHAEL [76] requires 60K A100 GPU days. As a result, cost-effective training strategies for linear Diffusion Transformers are still worth exploring.
In this work, we systematically study the diffusion Transformer implementation by pure linear attention, from both architectural design and training strategy. For architectural refinement, we draw on successful practices from multimodal understanding models and observe the free lunch effect of the linear attention. For training strategies, we focus on two cost-effective methods—weight inheritance and knowledge distillation—to improve the training process.
Our findings about architectural refinement and optimization strategies on efficient linear DiT can be summarized as the following practical guidelines: 1) employ Simplified Linear Attention in image understanding tasks; 2) use few (e.g., 2) heads in the linear attention for better performance but no latency increase; 3) initialize the weights of the linear diffusion Transformer from a well-trained self-attention based diffusion model; 4) load all weights except for the linear attention to achieve cost-effective training; 5) adopt hybrid distillation objective—distilling not only the predicted noise but also variances of the reverse diffusion process—to facilitate the student diffusion Transformer.
| (a) Comparison of subquadratic computational complexity designs | (b) Roadmap for improving LiT |
@@
These guidelines lead us to develop LiT, a diffusion Transformer using pure linear attention. LiT is trained in a highly cost-effective manner while maintaining a high-resolution-friendly property during inference, and can be deployed offline on a Windows 11 laptop. In class-conditional image generation on the ImageNet [13] 256256 benchmark, LiT-S/B/L (patch size of 2) with only 100K training steps outperforms corresponding 400K steps pre-trained DiT-S/B/L in terms of Fréchet Inception Distance (FID) [27]. For large-scale model on the ImageNet 256256 and 512512 benchmark, LiT-XL/2 competes comparably to DiT-XL/2 (2.32 vs. 2.27, 3.69 vs. 3.04 in FID) with only and of training steps (respectively for 256256 and 512512 settings). In challenging text-to-image generation, our LiT-0.6B can stably generate highly photorealistic images (Fig. 7) with only 40K training steps. As a result, we summarize the main contributions:
We introduce a linear diffusion Transformer, LiT, which optimizes the latency-capacity trade-off by using a few heads (e.g., 2) in linear attention.
We initialize LiT with weight inheritance to ensure a cost-effective training process, while achieving linear computational complexity that facilitates high-resolution generation scenarios during inference.
We propose a hybrid distillation approach, thereby successfully apply LiT to class-conditional image generation and text-to-image generation.
The pretrained LiT-0.6B model can be deployed on a Windows 11 laptop to generate high-resolution, photo-realistic images offline based on user prompts.
2 Related Work
Diffusion models [29, 67, 68] have proven effective in image generation, and can be implemented using Transformer architecture based on self-attention [71], e.g., DiT [53], PixArt- [8], PixArt- [7], and Lumina-T2X [18].
From an architectural perspective, recent work begins focusing on efficient diffusion models [69, 77, 34], based on Mamba [21, 11] and RWKV [54] to compete with DiT. Meanwhile, linear attention [38], as a simple linear complexity token mixer [83], is being applied in diffusion Transformers for image synthesis, e.g., Mediator [58], DiG [87] for class-conditional image generation, and Sana [75], LinFusion [46] for text-to-image generation. In summary, existing work trains diffusion Transformers with linear attention from scratch. In contrast, our LiT aims for cost-effective training by exploring linear attention design, weight inheritance and knowledge distillation. Moreover, our approach has been validated on both class-conditional image generation and text-to-image generation.
3 Preliminaries and Motivation
3.1 Preliminaries
@@
The computational complexities of softmax attention and linear attention are and , respectively. Detailed presentation is provided in Sec. 4.2.
| (a) Latency vs resolution | (b) GPU memory vs resolution | (c) Cosine similarity among heads |
3.2 Motivation
For scenarios requiring high-resolution generation (e.g., 10241024), self-attention is latency-intensive. In a common setup, using a latent diffusion Transformer [62] with an 8 downsampling VAE [39] encoder (as in DiT-XL/2 [53]), the input latent resolution is 128128, bringing a sequence length of 4096. Accordingly, we have . For higher resolutions, this ratio will increase.
Based on this analysis, self-attention is not well-suited for commercially valuable high-resolution generation tasks. Can we use cheap linear attention to replace self-attention in diffusion models? To this end, we aim to explore how to cost-effectively train a linear DiT with photorealistic generation quality by focusing on linear attention design, weight inheritance, and diffusion knowledge distillation paradigm. We discuss our experiments and conclusions in detail below.
4 Cost-Effectively Training of Linear DiT
In this section, we provide the LiT design trajectory. Given the limited research on linear DiT, we first build a suitable linear attention baseline for generative tasks, before investigating advanced training strategies. Our baseline shares the same macro architecture as DiT [53], which has recently proven effective for text-to-image generation [8, 7, 44], with the only difference being that we replace self-attention with linear attention that is computationally efficient in high-resolution image synthesis. Additionally, we employ latent diffusion [62] consistent with DiT. LiT generally follows DiT’s configuration (for S/B/L/XL), except the number of attention heads. All experiments are conducted on the class-conditional ImageNet [13] 256256 benchmark, with models trained for 400K iterations using a batch size of 256. We report the FID-50K [27], Inception Score [65] and Precision/Recall [40] metrics. The radar chart of the experimental results is shown in Fig. 2-(b). Detailed results can be found in Sec. D of the appendix.
4.1 Simplified Linear Attention in DiT
| DiT | Attention | FID-50K () | IS () |
| S/2 | softmax | 68.40 | - |
| S/2 | ReLU linear | 88.46 | 15.11 |
| S/2 | Simplified linear (ReLU) | 63.66 | 22.16 |
| S/2 | focused linear (ReLU) | 63.05 | 22.49 |
| S/2 | focused linear (GELU) | 70.83 | 19.41 |
| B/2 | softmax | 43.47 | - |
| B/2 | ReLU linear | 56.92 | 25.80 |
| B/2 | Simplified linear (ReLU) | 42.11 | 34.60 |
| B/2 | focused linear (ReLU) | 40.58 | 35.98 |
| B/2 | focused linear (GELU) | 58.86 | 24.23 |
Guideline 1: Simplified linear attention in image recognition is sufficient for DiT-based image generation.
Linear attention has been successfully applied in perceptual ViTs [3, 23, 24, 1] as a replacement for self-attention, which are used to extract high-level features for image classification. On the other hand, pioneering works apply them in generative models [58, 75, 46], where they are used to predict low-level noise for text-to-image synthesis. Inspired by the successful practices in ViTs [15], we now investigate how to simplify the linear attention design in LiT while effectively preserving the fidelity of generated images. We start with a common practice, i.e., ReLU linear attention [38], which uses the ReLU activation as a kernel function to approximate the non-linearity of the softmax function. For performance reference, we compare it to DiT, where any performance difference can be attributed to the impact of linear attention on generation quality. As shown in Tab. 1, compared to DiT, LiT-S/2 and B/2 with a trivial ReLU linear attention leads to an unacceptable performance drop, with FID decreasing by approximately 20/13 for the S/2 and B/2 models. The results suggest that linear attention commonly used in visual recognition have room for improvement in noise prediction tasks.
We then explore the following approaches: (1) simplified linear attention [22] (ReLU linear attention with a depthwise convolution (DWC) [33, 9]; (2) focused linear attention [23]; (3) replacing the ReLU kernel with GELU [26] function, commonly used in Transformers [71]. Each of these choices maintains linear complexity, preserving LiT’s advantage in computational efficiency. We use a relatively large convolution kernel (i.e., 5) to ensure a sufficiently large receptive field when predicting noise. Tab. 1 shows that adding DWC alone improves generation quality beyond softmax attention, with FID of 63.66 and 42.11 for S/2 and B/2. We attribute the result to the model’s tendency to focus on information from neighboring pixels when predicting noise for a given pixel, thus requiring the assistance of convolution. Meanwhile, we identify that the focused function has limited effectiveness, which we attribute to its design motivation to help linear attention focus on specific regions. This feature may suit classification models, but may not be necessary for noise prediction.
Remark 1. Given that a major motivation for using linear attention is to address latency issues with high-resolution images, we use simplified linear attention for simplicity and performance. Next, we focus on heads of the linear attention for a better latency-performance trade-off.
| (a) Results of small (S) and base (B) model | (b) Results of large (L) and Xlarge (XL) model |
4.2 Linear Attention with Few Heads
Theoretically, is independent of the number of heads, but in linear attention, has a negative correlation with it. Intuitively, using many heads can loose the computational stress. But on the contrary, we prefer to use few heads, based on the identified free lunch effect of linear attention. We illustrate the effect in Fig. 5, where we visualize the latency and theoretical GMACs of linear attention in DiT-S/B/L/XL models with different head numbers. We use one NVIDIA A100 GPU to generate 256256 resolution images with a batch size of 8, averaging the results over 30 experiments. The results show that decreasing the number of heads leads to a steady increase in theoretical GMACs but not the practical latency, which even shows a decrease. Accordingly, we summarize it as the free lunch effect of linear attention. We also observe this phenomenon on the standard NVIDIA V100 GPU, as reported in Sec. C of the appendix. As such, we argue that setting a small number of heads enables a high theoretical compute budget, which, according to scaling law [45, 37, 31], allows the model to reach a higher upper bound in generation performance. The argument can be empirically demonstrated in Tab. 2. For different model scales, equipping linear attention with a few heads (e.g., 2, 3, 4) outperforms the default settings in DiT. Instead, using an excessive number of heads (e.g., 96 for S/2 or 192 for B/2) seriously hinders the generation quality.
Besides, we also evaluate the necessity of using many heads in linear attention from a representation similarity lens. Specifically, we visualize the average cosine similarity among the attention maps of different heads in LiT-S/2 (with 6 heads). As depicted in Fig. 4-(c), for different denoising steps or layers, the attention across different heads is quite homogeneous, with the average cosine similarity reaching up to above 0.5. The observation suggests that a small number of heads dominate the major information in the linear attention, which positively supports the empirical generative results in Tab. 2.
@@ 4.2 Linear Attention with Few Heads
4.3 Inheriting Weights from Diffusion Transformer
LiT shares some structural components with DiT, allowing weights to be inherited from the pre-trained DiT. These weights contain rich knowledge related to noise prediction, which is expected to transfer to LiT in a cost-effective way. Therefore, we load the pre-trained DiT weights into LiT, except for the linear attention. This includes parameters related to FFN, adaptive layer norm (adaLN), positional encoding, and conditional embedding. To thoroughly investigate this, we conduct ablation experiments to explore how the degree of training on inherited weights affects the model’s generation performance.
Remark 3. We suppose a possible explanation for this phenomenon is the functional decoupling of different modules in diffusion Transformers, as explored in large language models for content generation [79, 73]. Basically, although DiT and LiT have different architectures, their shared components (e.g., FFN and adaLN) act quite similarly. As a result, the knowledge in the pre-trained parameters of these components can be transferred, and extensive pre-training may not hinder the optimization of the non-transfered parts.
@@ 4.3 Inheriting Weights from Diffusion Transformer
In LiT, some of the weights in linear attention overlap with those in self-attention of DiT, including the query, key, value, and output projection matrices. Despite the differences in computation paradigms, these weights can be directly loaded from DiT into the LiT without additional processing. However, whether this can accelerate its convergence remains an open question. We use a DiT-S/2 pre-trained for 600K iterations to conduct ablation experiments. Five different types of loading strategies are explored, including (1) query, key, and value projection matrices; (2) key and value projection matrices; (3) the value projection matrix; (4) the query projection matrix; and (5) the output projection matrix. Evaluation results are presented in Tab. 3. In comparison to the baseline without loading self-attention weights, none of the explored strategies shows better generative performance. The phenomenon can be attributed to differences in computation paradigms. Specifically, linear attention directly computes the product of key and value matrices, whereas self-attention does not. Consequently, weights related to key and value from self-attention offer limited benefit to linear attention. Clearly, as presented in Tab. 3, inheriting the weights of the key and value projector leads to a larger degradation (FID of 55.29 and 55.07).

### v2

# LiT: Delving into a Simple Linear Diffusion Transformer for Image Generation

@@ Abstract
In this paper, we investigate how to convert a pre-trained Diffusion Transformer (DiT) into a linear DiT, as its simplicity, parallelism, and efficiency for image generation. Through detailed exploration, we offer a suite of ready-to-use solutions, ranging from linear attention design to optimization strategies. Our core contributions include 5 practical guidelines: 1) Applying depth-wise convolution within simple linear attention is sufficient for image generation. 2) Using fewer heads in linear attention provides a free-lunch performance boost without increasing latency. 3) Inheriting weights from a fully converged, pre-trained DiT. 4) Loading all parameters except those related to linear attention. 5) Hybrid knowledge distillation: using a pre-trained teacher DiT to help the training of the student linear DiT, supervising not only the predicted noise but also the variance of the reverse diffusion process. These guidelines lead to our proposed Linear Diffusion Transformer (LiT), which serves as a safe and efficient alternative baseline for DiT with pure linear attention. In class-conditional 256256 and 512512 ImageNet generation, LiT can be quickly adapted from DiT using only and of DiT’s training steps, respectively, while achieving comparable performance. LiT also rivals methods based on Mamba or Gated Linear Attention. Moreover, the same guidelines generalize to text-to-image generation: LiT can be swiftly converted from PixArt- to generate high-quality images, maintaining comparable GenEval scores.
@@ 1 Introduction
Diffusion models [30, 68] based on Transformers [7, 16, 8] show strong commercial potential in image generation but face latency and GPU memory challenges in high-resolution tasks. Their main self-attention modules exhibit quadratic complexity with respect to sequence length, making deployment on resource-limited edge devices less feasible, as illustrated in Fig. 4 and Fig. 1 in Appendix Sec. B, highlighting the need for sub-quadratic attention modules.
Linear attention can significantly accelerate image generation compared to self-attention at the same resolution. For example, at 2048px, it is nearly 9 faster (Fig. 4-(a), (b)). Additionally, it substantially reduces GPU memory usage for high-resolution image generation. Replacing self-attention with linear attention in DiT-S/2 reduces memory consumption from 14GB to 4GB at 2048px. In fact, linear attention has been extensively validated in visual recognition tasks (e.g., image classification and segmentation) [3, 24, 1]. In these methods, the linear attention is expected to generate sharp distributions that effectively capture semantic information [3, 24]. However, in the context of denoising diffusion models, linear attention may require enhanced locality to better model the distribution of random noise. Consequently, exploring the use of linear attention in image generation remains a promising research direction.
Several efforts have explored the use of efficient attention for image generation. For instance, DiG [88] and DiM [71] reference Gated Linear Attention (GLA) [80] and State Space Models (SSM) [21, 22] to implement diffusion models, respectively. SANA [78] takes a deep compression autoencoder [9], and replace the original MLP-FFN with Mix-FFN [3]. Attention Mediators [59] applies a set of additional tokens in linear attention. CLEAR [46] adopts a circular window within the attention.
In this paper, we systematically investigate the problem of converting a pre-trained DiT to a linear DiT for image generation while minimizing performance loss. We decompose this problem into two parts: first, what type of linear attention architecture is suitable for image generation? Second, what training strategies are effective for optimizing a linear DiT? To address these two questions, we conducted extensive experiments on Linear DiT from both architecture design and training strategy perspectives. Based on the results, our findings can be summarized into five practical guidelines for Linear DiT design and optimization: 1) simply adding a depth-wise convolution in linear attention is sufficient for image generation; 2) a free lunch: few (e.g., 2) heads in linear attention improves performance without increasing latency; 3) initialize the weights of the linear DiT from a well-trained DiT; 4) load all weights except for the linear attention to achieve cost-effective training; 5) hybrid distillation objective—distilling not only the predicted noise but also variances of the reverse diffusion process—to facilitate the student linear DiT. Following such guidelines (Fig. 3), the result diffusion model, dubbed Linear Diffusion Transformer (LiT), can adopt pure simple attention design, achieving linear computational complexity while retaining the capability of full attention. Meanwhile, LiT is trained in a highly cost-effective manner as it is derived from a pre-trained DiT, leveraging existing knowledge and reducing the need for extensive training from scratch.
@@
LiT is an affirmative answer to how to efficiently and safely obtain a linear DiT from a pretrained DiT. To validate its effectiveness, we conduct experiments on representative diffusion Transformers, i.e., DiT (for ImageNet [14] class-conditional image generation) and PixArt- [7] (for text-to-image generation). In ImageNet [14] 256256 benchmark, LiT-S/B/L (patch size of 2) with only 100K training steps outperforms corresponding 400K steps pre-trained DiT-S/B/L in terms of Fréchet Inception Distance (FID) [28]. For large-scale model on the ImageNet 256256 and 512512 benchmark, LiT-XL/2 competes comparably to DiT-XL/2 (2.32 vs. 2.27, 3.69 vs. 3.04 in FID) with only and of training steps (respectively for 256256 and 512512 settings). In text-to-image generation, converted from PixArt-, LiT-0.6B shares the same macro/micro-level architecture, ensuring alignment with the PixArt- framework while elegantly replacing all self-attention with efficient linear attention. LiT-0.6B can stably generate highly photorealistic images (Fig. 2). To verify the effectiveness of linear attention, we deploy LiT-0.6B offline on a Windows 11 laptop (Fig. 11), enabling 1K-resolution image synthesis based on human instructions.
2 Preliminaries
@@
| DiT | Attention | FID-50K () | IS () |
| S/2 | Softmax | 68.40 | - |
| S/2 | ReLU Linear Baseline | 88.46 | 15.11 |
| S/2 | + Depth-wise Conv. (ReLU) | 63.66 | 22.16 |
| S/2 | + Focused Linear (ReLU) | 63.05 | 22.49 |
| S/2 | + Focused Linear (GELU) | 70.83 | 19.41 |
| B/2 | softmax | 43.47 | - |
| B/2 | ReLU Linear Baseline | 56.92 | 25.80 |
| B/2 | + Depth-wise Conv. (ReLU) | 42.11 | 34.60 |
| B/2 | + Focused Linear (ReLU) | 40.58 | 35.98 |
| B/2 | + Focused Linear (GELU) | 58.86 | 24.23 |
@@
The computational complexities of softmax attention and linear attention are and , respectively.
Despite its efficiency, the suitability of linear attention for image generation remains under-explored. Most existing works, such as SANA [78], Mediators [59], and CLEAR [46], introduce carefully designed modifications to DiT, compromising its simplicity. Motivated by this, we explore how to safely and efficiently convert a pretrained DiT into a linear DiT without relying on complex architectures or sacrificing performance.
3 Exploring LiT: A Roadmap
We explore the architectural design and optimization strategies of our LiT based on DiT [54] on class-conditional ImageNet [14] 256256 benchmark, as it is relatively simple, thereby highlighting the ablation effect of attention itself. We train models for 400K iterations using a batch size of 256, and report the FID-50K [28], Inception Score [66] and Precision/Recall [40] metrics. The radar chart of the experimental results is shown in Fig. 1. Detailed results can be found in Sec. E of the Appendix.
3.1 Architectural Guidelines for Linear DiT
| (a) Latency | (b) GPU Memory |
@@ 3.1 Architectural Guidelines for Linear DiT
Guideline 1: Simply adding a depth-wise convolution in linear attention is sufficient for DiT-based image generation.
In high-level visual recognition tasks (e.g., image classification, instance segmentation, etc.), linear attention has been reported to produce a “smooth” attention distribution compared to the “sharp” distribution of softmax attention [3, 24]. To address the issue, Flatten Transformer [24] introduces a focused function applied after the ReLU activation as a remedy. Meanwhile, linear attention has been shown to exhibit limited feature diversity due to its low-rank attention matrix. Accordingly, only a single depth-wise convolution layer, applied on the value matrix, can raise the upper bound of the equivalent rank, effectively mitigating the low-rank issue [24]. Based on these observations, it is reasonable to validate similar conclusions in image generation. Compared to the full attention baseline, LiT-S/2 and B/2 with ReLU linear attention exhibit an unacceptable performance drop (Tab. 1), with FID increasing by approximately 20/13 for the S/2 and B/2 models, respectively. We then explore the following approaches: (1) adding a depth-wise convolution (DWC) [34, 10] on the value matrix111In this setup, we optimize the linear attention implementation (e.g., by replacing einsum with the “@” operator for matrix operations and removing the softplus function). We also test a version without these optimizations for comparison, which yielded results of 63.13 (S/2) and 42.13 (B/2), showing only slight differences from the results in Tab. 1: 63.66 (S/2) and 42.11 (B/2). Therefore, we retain the optimizations.; (2) adding focused function [24]; (3) replacing the ReLU kernel with GELU [27] function, commonly used in Transformers [73]. Each of these choices maintains linear complexity, preserving LiT’s advantage in computational efficiency. We use a relatively large convolution kernel (i.e., 5) to ensure a sufficiently large receptive field when predicting noise.
As shown in Tab. 1, incorporating DWC alone enhances generation quality beyond softmax attention, achieving FID scores of 63.66 and 42.11 for S/2 and B/2, respectively. We attribute this improvement to the model’s tendency to rely on information from neighboring pixels when predicting noise for a given pixel, necessitating the assistance of convolution. Meanwhile, we identify that the focused function has limited effectiveness, which we attribute to its design motivation to help linear attention focus on specific regions. This feature may suit classification models, but may not be necessary for noise prediction.
Remark 1. We note that our findings on linear attention in visual generation align with those of the SLAB [23] in visual recognition, where simple DWC leads to improvement. We attribute this to linear attention’s lack of locality modeling capacity compared to full attention, which DWC serves as a remedy. We will adopt this design afterwards.
@@
3.2 Linear Attention with Few Heads
| (a) Small/Base | (b) Large/XLarge |
Theoretically, is independent of the number of heads, but in linear attention, has a negative correlation with it. Intuitively, using many heads can loose the computational stress. But on the contrary, we prefer to use few heads, based on the identified free lunch effect of linear attention. We illustrate the effect in Fig. 5, where we visualize the latency and theoretical GMACs of linear attention in DiT-S/B/L/XL models with different head numbers. We use one NVIDIA A100 GPU to generate 256256 resolution images with a batch size of 8, averaging the results over 30 experiments. The results show that decreasing the number of heads leads to a steady increase in theoretical GMACs but not the practical latency, which even shows a decrease. Accordingly, we summarize it as the free lunch effect of linear attention. We also observe this phenomenon on NVIDIA V100 GPU, as reported in Fig. 3 in Appendix Sec. C. As such, we argue that setting a small number of heads enables a high theoretical compute budget, which, according to scaling law [44, 38, 32], allows the model to reach a higher upper bound in generation performance. The argument can be empirically demonstrated in Tab. 2. For different model scales, equipping linear attention with a few heads (e.g., 2, 3, 4) outperforms the default settings in DiT. Instead, using an excessive number of heads (e.g., 96 for S/2 or 192 for B/2) seriously hinders the generation quality.
Besides, we also evaluate the necessity of using many heads in linear attention from a representation similarity lens. Specifically, we visualize the average cosine similarity among the attention maps of different heads in LiT-S/2 (with 6 heads). As depicted in Fig. 4 in Appendix Sec. D, for different denoising steps or layers, the attention across different heads is quite homogeneous, with the average cosine similarity reaching up to above 0.5. The observation suggests that a small number of heads dominate the major information in the linear attention, which positively supports the empirical generative results in Tab. 2.
3.3 Inheriting Weights from DiT
LiT can be converted from a pre-trained DiT, whose weights contain rich knowledge of noise prediction, enabling a cost-effective transfer to LiT. Thus, we load the pre-trained DiT weights into LiT, except for the linear attention. DiT can be pretrained to varying convergence levels. Given the architectural differences between DiT and LiT, whether converged DiT weights serve as optimal initialization for LiT remains an open question. To this end, we conduct ablation experiments to explore how the degree of training on inherited weights affects the model’s generation performance.
Remark 3. We suppose a possible explanation for this phenomenon is the functional decoupling of different modules in DiTs, as explored in large language models [81, 75]. Basically, although DiT and LiT have different architectures, their shared components (e.g., FFN and adaLN) act quite similarly. As a result, knowledge in these components can be transferred, and extensive pre-training may not hinder the optimization of the non-transfered parts.